OpenStudy (anonymous):

Factorise 4a³b +8a²b² - 4ab² - 2a²b

OpenStudy (whpalmer4):

Do you see any common factors?

OpenStudy (anonymous):

2ab

OpenStudy (whpalmer4):

Okay, what do you get after you've factored that out?

OpenStudy (anonymous):

2ab (2a² + 4ab - 2b - a)

OpenStudy (whpalmer4):

\[4a^3b+8a^2b^2-4ab^2-2a^2b =2ab(2a^2+4ab-2b-a)\] Yes. Can you factor the expression in the parentheses?

OpenStudy (whpalmer4):

Hint: factor by grouping

OpenStudy (whpalmer4):

Taking just the part inside the parentheses for now: \[2a^2+4ab-2b-a = (2a^2+4ab) - (2b+a) = 2a(a+2b)-(2b+a) \]\[= (2a-1)(a+2b)\] So our entire thing factored is \[2ab(2a-1)(a+2b)\]

OpenStudy (anonymous):

Thankyou I get it now